A sample of 2,4-hexadiene in methanol in a 2.5 cm cell shows an
absorbance of 0.65 at 240.0 nm. What is the concentration of the
diene? The molar absorption at this wavelength is 12100 dm3 mol1
cm1
.
1
Expert's answer
2019-10-28T07:50:13-0400

Absorbance

"=A=\epsilon\times c\times l"

Absorbance of the solution at 240 nm = 0.65

Molar concentration of diene = "c"

Length of the cell = 2.5 cm

The molar absorption at this wavelength

"=12100"

Now,


"c=\frac{A}{\epsilon \times l}=\frac{0.65}{1200\times 2.5}=2.14\times 10^{-5}\ moles\ l^{-1}"